A Car travelling at a constant speed of 45.0m/s passes a highway patrol(HP) hidden behind a billboard. A second after the car passes the billboard the HP sets out to catch it, accelerating at a constant 3.00s. How long does it take to overtake the car?


1
Expert's answer
2021-09-14T09:36:06-0400

Given:

"v=45.0\:\rm m/s"

"a=3.00\:\rm m/s^2"


The equations of motion of the cars

"d_1=vt,\:\quad d_2=at^2/2"

We get

"d_1=d_2"

"vt=at^2\/2"

"t=2v\/a"

"t=2*45.0/3.00=30\:\rm s"